OpenStudy (wesdg1978):

My answer isn't exactly matching the signs of the given answer and I know this may sound like a stupid question but I've confused myself. Does xy+x+y=x^2y^2 become xy+x+y-x^2y^2=0 OR xy+x+y+x^2y^2=0

OpenStudy (amistre64):

xy+x+y=x^2y^2 is this your initial function?

OpenStudy (wesdg1978):

Yes

OpenStudy (amistre64):

and what are you trying to do with it?

OpenStudy (wesdg1978):

Well I did implicit differentiation and got the answer and my answer is the same as one of the four options, except for the sign.

OpenStudy (amistre64):

xy + x + y = x^2y^2 xy: x'y + xy' x: x' y: y' x^2y^2: x^2' y^2 + x^2 y^2' : 2xx' y^2 + x^2 2yy' assuming this is "with respect to x"; x' = dx/dx = 1 y + xy' + 1 + y' = 2x y^2 + x^2 2yy' collect the y's to factor out xy' + y' - x^2 2yy' = 2x y^2 - y - 1 y' (x + 1 - 2x^2 y) = 2x y^2 - y - 1 and divide \[y' = \frac{2x y^2 - y - 1}{x + 1 - 2x^2 y}\] simplify as wanted

OpenStudy (wesdg1978):

Yeah, that's the answer they give, but this is what I got: